OpenStudy (anonymous):

Sue buys 5 notebooks and 3 binders for 21.56. Tim buys 4 notebooks and 2 binders for 15.80. how much is each item?

OpenStudy (anonymous):

There are two values we do not know - the cost of a notebook and the cost of a binder Let x be the cost of a notebook. Let y be the cost of a binder. 5*(cost of notebook) + 3*(cost of binder) = 21.56 4*(cost of notebook) + 2 * (cost of binder) = 15.80 Replacing with x and y 5x + 3y = 21.56 4x + 2y = 15.80 Ok so far?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):